# Integrating GNN and Neural ODEs for Estimating Two-Body Interactions in Mixed-Species Collective Motion

### v1

@@ 5 Experiments
In this section, we describe two numerical experiments to demonstrate the efficacy of the proposed method.
@@ 5.1 Underdamped Brownian Motion with Harmonic Interaction
Here, $k=1\times 10^{-3}$, $r_{c}=1$, $\sigma=1\times 10^{-3}$, and $\gamma$ was set to $1\times 10^{-2}$ or $1\times 10^{-1}$. The initial states $x^{i}(0)$ were uniformly sampled from $\mathcal{X}_{0}=[0,5]^{2}\times[0,1\times 10^{-3}]$, and simulations were performed to obtain $Z_{\text{sim}}$.
@@
$F^{(1)}_{\text{NN},v}$ and $F^{(2)}_{\text{NN},v}$ were each independent neural networks. We evaluated the predictive accuracy for position and velocity as follows, using $Z^{m}_{\text{NN}}(\theta;t_{0}+\tau)$ predicted for $\tau=10$ and the simulation data $Z^{m}_{\text{sim}}(t_{0})$:
@@
The accuracy of the estimation results is quantified in Supplemental Table S1. For each particle or pair in the dataset $Z_{\text{sim}}$, we calculated the forces $F^{(1)}_{v}$ and $F^{(2)}_{v}$ and computed the Mean Squared Error (MSE) and Mean Absolute Error (MAE). These errors were then normalized by the L1 or L2 norm of $F^{(1)}_{\text{sim},v}$ and $F^{(2)}_{\text{sim},v}$ respectively, to provide dimensionless measures of accuracy. Due to the extensive computation time required for estimation, comprehensive statistics could not be gathered. Instead, we present the results of all trials to illustrate the trends in estimation accuracy. As observed, when $M_{\text{tra}}=3$ with a low friction constant $\gamma=1\times 10^{-2}$, significant estimation errors occurred. Conversely, the estimation errors for the interaction force $F^{(2)}_{v}$ were largely unaffected by $\gamma$, suggesting that the estimation of interaction and friction are somewhat independent. However, when $M_{\text{tra}}=270$, a slight improvement in the accuracy of interaction estimates was noted compared to the $M_{\text{tra}}=3$ scenario. This observation indicates that not only friction but also the accuracy of interaction estimates depend on the number of data points used.
Furthermore, to verify how well these estimates fit the simulation data, we sampled random initial values $Z(0)$ in a similar manner to training data creation, conducted simulations, and visualized the results $S_{\text{NN}}(\theta^{*})(Z(0),t)$ (Figure 1(I-K); Supplemental Movies S3-5). In all cases, the estimates were confirmed to adequately replicate the simulation data for all cases, even with $\gamma=1\times 10^{-2}$ and $M_{\text{tra}}=3$. The ability to replicate simulation data suggests that friction had almost no effect in this case. However, this implies that the proposed method may not adequately estimate very weak effects.
@@ 5.2 Mixed Species Collective Motion with Overdamped Self-propulsion
Figure 2: Snapshots of the simulation results in the mixed-species model. Panel (A) displays the case with parameter set (i), and panel (B) with parameter set (ii). Each individual is represented by an arrow located at their position and directed toward their polarity. Black arrows indicate individuals with $c^{i}=0$, while red arrows represent those with $c^{i}=1$. (C-J) The functions estimated from the data with the mixed-species model plotted against the true values. The rows display different cases: (C-F) with parameter set (i), and (G-J) with parameter set (ii). A gray line serves as a guide indicating where the estimated values equal the true values. (K-L) Snapshots of the simulation results in the model estimated from data with the mixed-species model. (K) The case with parameter set (i), and (L) with parameter set (ii).
Next, to test the proposed method in complex systems involving interactions among multiple species, we conducted simulations using a more complex model that emulates real collective movements. In this model, each of the $N=400$ individuals has position $r^{i}(t)\in[0,L]^{2}$, orientation $\phi^{i}(t)\in[0,2\pi]$, and species type $c^{i}\in\{0,1\}$. Both $r^{i}$ and $\phi^{i}$ are subject to periodic boundary conditions, ensuring continuous and consistent movement dynamics across the defined space. This defines $x^{i}(t)=(r^{i}(t),\phi^{i}(t))$ and $y^{i}=(c^{i})$, constituting $z^{i}(t)=(x^{i}(t),y^{i})$. The motion equations for individual $i$ are described as follows:
@@
$r^{ij}$ $=$ $r^{j}-r^{i}\in[-L/2,L/2]^{2}.$ (22)
with self-propulsion speed $v_{0}\geq 0$, the strength of the excluded volume interaction $\beta\geq 0$, the strengths of contact following and chemotaxis $\alpha_{\text{CF}},\alpha_{\text{Ch}}\geq 0$ , respectively, the diffusion length of the chemoattractant $\kappa\geq 0$, the noise strength $\sigma\geq 0$, and the modified Bessel function of the second kind $K_{1}$. The model was obtained by introducing a chemotaxis term into a preexisting model [14], to make it more appropriate for cellular slime molds. The chemotaxis term assumes rapid diffusion of chemotactic substances secreted by each individual [20]. We varied the sensitivity of contact following and chemotaxis depending on the species type, thus modeling a system where different species interact. The interaction terms were adapted into our framework as follows:
@@
Using this data, we applied our method to estimate $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$, modeling $F^{(1)}_{\text{NN}}$ and $F^{(2)}_{\text{NN}}$ as follows:
@@
To minimizing this loss function, we tried the same hyperparameters as those used for the harmonic interaction model, and found we optimized $\theta$ to estimate $F^{(1)}_{\text{NN}}$ and $F^{(2)}_{\text{NN}}$ (Figure 2(C-J)). All experiments were conducted with $M_{\text{tra}}=3,M_{\text{val}}=3$. After 20,000 epochs (25 days), $F^{(1)}_{\text{NN},r},F^{(1)}_{\text{NN},\phi},F^{(2)}_{\text{NN},r},F^{(2)}_{\text{NN},\phi}$ were found to approximate $F^{(1)}_{\text{sim},r},F^{(1)}_{\text{sim},\phi},F^{(2)}_{\text{sim},r},F^{(2)}_{\text{sim},\phi}$ closely (Figure 2(C-J)). To further verify the fit of our estimation results with the simulation data, we sampled random initial values $Z(0)$ similarly to the training data creation process and conducted simulations to visualize the outcomes $S_{\text{NN}}(\theta^{*})(Z(0),t)$ (Figure 2(K-L); Supplemental Movies S8-9). The results confirmed that all estimates adequately reproduced the simulation data.
@@
To further test the predictive power of the model trained on the mixed state, we performed simulations where all individuals were of the same type. For initial conditions $Z_{c}(0)$ where all $N=400$ individuals shared the same type $c^{i}=c$, either $c=0$ or $c=1$, we performed simulations $S_{\text{sim}}(Z_{c}(0),t)$ and $S_{\text{NN}}(\theta^{*})(Z_{c}(0),t)$, visualizing each case (Supplemental Figure S1). For parameter set (i) with $c^{i}=0$, both $S_{\text{sim}}(Z_{0}(0),t)$ and $S_{\text{NN}}(\theta^{*})(Z_{0}(0),t)$ showed the formation of two clusters that exhibited rotational movements (Supplemental Figure S1(A-B)). Conversely, for parameter set (i) with $c^{i}=1$ and parameter set (ii) with $c^{i}=0,1$, both $S_{\text{sim}}(Z_{0}(0),t)$ and $S_{\text{NN}}(\theta^{*})(Z_{0}(0),t)$ formed centipede-like clusters that moved translationally in alignment with their local direction (Supplemental Figure S1(C-H)). Finally, as a negative control, we conducted simulations where both types of individuals shared the same parameters, as in parameter sets (iii)-(v), and performed estimations using the neural network for each case (Supplemental Figure S2). In all instances, it was confirmed that $F^{(1)}_{\text{NN},r},F^{(1)}_{\text{NN},\phi},F^{(2)}_{\text{NN},r},F^{(2)}_{\text{NN},\phi}$ closely approximated $F^{(1)}_{\text{sim},r},F^{(1)}_{\text{sim},\phi},F^{(2)}_{\text{sim},r},F^{(2)}_{\text{sim},\phi}$. Notably, these estimation results showed minimal dependency on $c^{i}$ in $F^{(1)}_{\text{NN}}(z^{i};\theta^{*}),F^{(2)}_{\text{NN}}(z^{i},z^{j};\theta^{*})$. This outcome suggests that the species dependency of interactions estimated by the proposed method in mixed assemblies is based on actual data, rather than being an artifact of the methodology.
@@
As detailed in Supplemental Table S3, we quantified the estimation accuracy for each trial, similar to the methods described in Section 5.1. To note, since $F^{(1)}_{\text{sim},\phi}=0$, the MSE and MAE for $F^{(1)}_{\phi}$ are not normalized. Our results demonstrate that $F^{(1)}_{r}$ is estimated with high accuracy across all conditions. Also, the estimation of $F^{(1)}_{\phi}$ is maintained at sufficiently small magnitudes relative to the order of magnitude ($10^{0}$) typically seen with $F^{(2)}_{\phi}$ (Figure 2(F,J)). Regarding the interactions, both $F^{(2)}_{r}$ and $F^{(2)}_{\phi}$ exhibited somewhat higher errors, approximately of the order of $10^{-1}$. These errors likely stem from the high dimensionality of the inputs to the $F^{(2)}_{\text{NN}}$, which may prevent the neural network from achieving sufficient learning, or from the training process neglecting the noise.
@@ 6 Conclusion
In this study, we proposed a novel method for estimating interactions among individuals within models of collective motion. This method combines dynamic graph neural networks with Neural Ordinary Differential Equations (Neural ODEs) to estimate interactions among individuals. We demonstrated that this method could estimate interactions in both simple models and complex mixed-species collective motion models. In models with simple harmonic interactions, the proposed method was able to identify crucial interactions effectively. Particularly, in the complex mixed-species collective motion model, it was estimated that individuals of different species exhibited distinct interactions, demonstrating the efficacy of the proposed method. One limitation of this approach is the extensive time required for estimation. Moreover, the method currently estimates deterministic motion equations and only considers pairwise interactions, thereby not accommodating interactions among three or more bodies or the effects of noise. Future research is expected to extend this approach to estimate more general interactions and develop methods for stochastic motion equations.
@@ Acknowledgements
This research was supported by JSPS KAKENHI Grant Numbers JP22H05673 awarded to MU, JP22H04841 and JP22K14012 awarded to SKS, JP19H05799 awarded to TJK, and JP19H05801 and JP19KK0282 awarded to SS. Additionally, support was provided by JST CREST Grant Numbers JPMJCR2011 to TJK and JPMJCR1923 to SS. We acknowledge the use of ChatGPT and GitHub Copilot for their assistance in swiftly articulating ideas and refining them into clear, comprehensible English text. The content was subsequently reviewed and finalized by the authors.

### v2

@@ 5 Experiments
Here, we demonstrate the efficacy of the proposed method in two numerical experiments.
@@ 5.1 Underdamped Brownian Motion with Harmonic Interaction
Here, $k=1\times 10^{-3}$, $r_{c}=1$, $\sigma=1\times 10^{-3}$, and $\gamma$ was set to $1\times 10^{-2}$ or $1\times 10^{-1}$. The initial states $x^{i}(0)$ were uniformly sampled from $\mathcal{X}_{0}=[0,5]^{2}\times[0,1\times 10^{-3}]^{2}$, and simulations were performed to obtain $Z_{\text{sim}}$.
@@
$F^{(1)}_{\text{NN},v}$ and $F^{(2)}_{\text{NN},v}$ were each independent neural networks as described in section 4.5. We evaluated the predictive accuracy for position and velocity as follows, using $Z^{m}_{\text{NN}}(\theta;t_{0}+\tau)$ predicted for $\tau=10$ and the simulation data $Z^{m}_{\text{sim}}(t_{0})$:
@@
The accuracy of the estimation results is quantified in Supplemental Table S1. For each particle or pair in the dataset $Z_{\text{sim}}$, we calculated the forces $F^{(1)}_{v}$ and $F^{(2)}_{v}$ and computed the Mean Squared Error (MSE) and Mean Absolute Error (MAE). These errors were then normalized by the L1 or L2 norm of $F^{(1)}_{\text{sim},v}$ and $F^{(2)}_{\text{sim},v}$ respectively, to provide dimensionless measures of accuracy. Due to the extensive computation time required for estimation, comprehensive statistics could not be gathered. Instead, we present the results of all trials to illustrate the trends in estimation accuracy. As observed, when $M_{\text{tra}}=3$ with a low friction constant $\gamma=1\times 10^{-2}$, significant estimation errors occurred. Conversely, the estimation errors for the interaction force $F^{(2)}_{v}$ were largely unaffected by $\gamma$, suggesting that the estimation of interaction and friction are somewhat independent. However, when $M_{\text{tra}}=270$, a slight improvement in the accuracy of interaction estimates was noted compared to the $M_{\text{tra}}=3$ scenario. This observation indicates that the accuracy of estimation depends on the number of data points, for the interaction as well as the friction.
Furthermore, to verify how well these estimates fit the training data, we conducted simulations from random initial values $Z(0)$ sampled in a similar manner to training data creation, and visualized the results $S_{\text{NN}}(\theta^{*})(Z(0),t)$ (Figure 2(I-K); Supplemental Movies S3-5). In all cases, the estimates were confirmed to adequately replicate the training data for all cases, even with $\gamma=1\times 10^{-2}$ and $M_{\text{tra}}=3$. The ability to replicate training data suggests that friction had almost no effect in this case. However, this implies that the proposed method may not adequately estimate very weak effects.
@@ 5.2 Mixed Species Collective Motion with Overdamped Self-propulsion
Figure 3: Snapshots of the simulation results in the mixed-species model. Panel (A) displays a representative snapshot of the training data for the case with parameter set (i), and panel (B) with parameter set (ii) (see Supplemental Table S2). Each individual is represented by an arrow located at their position and directed toward their polarity. Black arrows indicate individuals with $c^{i}=0$, while red arrows represent those with $c^{i}=1$. (C-J) The functions estimated from the data with the mixed-species model plotted against the true values. The rows display different cases: (C-F) with parameter set (i), and (G-J) with parameter set (ii). A gray line serves as a guide indicating where the estimated values equal the true values. (K-L) Snapshots of the simulation results in the model estimated from data with the mixed-species model. (K) The case with parameter set (i), and (L) with parameter set (ii).
Next, to test the proposed method in complex systems involving interactions among multiple species, we conducted simulations using a more complex model that emulates real collective movements.
Training data:
In this model, each of the $N=400$ individuals has position $r^{i}(t)\in[0,L]^{2}$, orientation $\phi^{i}(t)\in[0,2\pi]$, and species type $c^{i}\in\{0,1\}$. Both $r^{i}$ and $\phi^{i}$ are subject to periodic boundary conditions, ensuring continuous and consistent movement dynamics across the defined space. This defines $x^{i}(t)=(r^{i}(t),\phi^{i}(t))$ and $y^{i}=(c^{i})$, constituting $z^{i}(t)=(x^{i}(t),y^{i})$. The motion equations for individual $i$ are described as follows:
@@
$r^{ij}$ $=$ $r^{j}-r^{i}\in[-L/2,L/2]^{2},$ (22)
with self-propulsion speed $v_{0}\geq 0$, the strength of the excluded volume interaction $\beta\geq 0$, the strengths of contact following and chemotaxis $\alpha_{\text{CF}},\alpha_{\text{Ch}}\geq 0$ , respectively, the diffusion length of the chemoattractant $\kappa\geq 0$, the noise strength $\sigma\geq 0$, and the modified Bessel function of the second kind $K_{1}$. The model was obtained by introducing a chemotaxis term into a preexisting model [18], to make it more appropriate for cellular slime molds. The chemotaxis term assumes rapid diffusion of chemotactic substances secreted by each individual [24]. We varied the sensitivity of contact following and chemotaxis depending on the species type, thus modeling a system where different species interact non-reciprocally. The interaction terms were adapted into our framework as follows:
@@
Learning:
In order to estimate $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$, we modeled $F^{(1)}_{\text{NN}}$ and $F^{(2)}_{\text{NN}}$ as follows:
@@
To minimizing this loss function, we tried the same hyperparameters as those used for the harmonic interaction model, and found we optimized $\theta$ to estimate $F^{(1)}_{\text{NN}}$ and $F^{(2)}_{\text{NN}}$ (Figure 3(C-J)). All experiments were conducted with $M_{\text{tra}}=3,M_{\text{val}}=3$. After 20,000 epochs (25 days), $F^{(1)}_{\text{NN},r},F^{(1)}_{\text{NN},\phi},F^{(2)}_{\text{NN},r},F^{(2)}_{\text{NN},\phi}$ were found to approximate $F^{(1)}_{\text{sim},r},F^{(1)}_{\text{sim},\phi},F^{(2)}_{\text{sim},r},F^{(2)}_{\text{sim},\phi}$ closely (Figure 3(C-J)). To further verify the fit of our estimation results with the training data, we sampled random initial values $Z(0)$ similarly to the training data creation process and conducted simulations to visualize the outcomes $S_{\text{NN}}(\theta^{*})(Z(0),t)$ (Figure 3(K-L); Supplemental Movies S8-9). The results confirmed that all estimates adequately reproduced the training data.
@@
To further test the predictive power of the model trained on the mixed state, we performed simulations where all individuals were of the same type. For initial conditions $Z_{c}(0)$ where all $N=400$ individuals shared the same type $c^{i}=c$, either $c=0$ or $c=1$, we performed simulations using the aforementioned model $S_{\text{sim}}$ and the neural network $S_{\text{NN}}(\theta^{*})$ trained above (Supplemental Figure S1). For parameter set (i) with $c=0$, both $S_{\text{sim}}(Z_{0}(0),t)$ and $S_{\text{NN}}(\theta^{*})(Z_{0}(0),t)$ showed the formation of two clusters that exhibited rotational movements (Supplemental Figure S1(A-B)). Conversely, for parameter set (i) with $c=1$ and parameter set (ii) with $c=0,1$, both $S_{\text{sim}}(Z_{0}(0),t)$ and $S_{\text{NN}}(\theta^{*})(Z_{0}(0),t)$ formed centipede-like clusters that moved translationally in alignment with their local direction (Supplemental Figure S1(C-H)). These results suggest that the interaction of each type was approximated accurately enough to explain the dynamics of training data.
As a negative control, we conducted simulations where both types of individuals shared the same parameters, as in parameter sets (iii)-(v), and trained the neural networks for each case (Supplemental Figure S2). In all instances, it was confirmed that $F^{(1)}_{\text{NN},r},F^{(1)}_{\text{NN},\phi},F^{(2)}_{\text{NN},r},F^{(2)}_{\text{NN},\phi}$ closely approximated $F^{(1)}_{\text{sim},r},F^{(1)}_{\text{sim},\phi},F^{(2)}_{\text{sim},r},F^{(2)}_{\text{sim},\phi}$. Notably, these estimation results showed minimal dependency on $c^{i}$ in $F^{(1)}_{\text{NN}}(z^{i};\theta^{*}),F^{(2)}_{\text{NN}}(z^{i},z^{j};\theta^{*})$. This outcome suggests that the species dependency of interactions estimated by the proposed method in mixed assemblies is based on actual data, rather than being an artifact of the methodology.
@@
Finally, to summarize the results above, we quantified the estimation accuracy for each trial, similar to the methods described in section 5.1 (Supplemental Table S3). To note, since $F^{(1)}_{\text{sim},\phi}=0$, the MSE and MAE for $F^{(1)}_{\phi}$ are not normalized. Our results demonstrate that $F^{(1)}_{r}$ is estimated with high accuracy across all conditions. Also, the estimation of $F^{(1)}_{\phi}$ is maintained at sufficiently small magnitudes relative to the order of magnitude ($10^{0}$) typically seen with $F^{(2)}_{\phi}$ (Figure 3(F,J)). Regarding the interactions, both $F^{(2)}_{r}$ and $F^{(2)}_{\phi}$ exhibited somewhat higher errors, approximately of the order of $10^{-1}$. These errors likely stem from the high dimensionality of the inputs to the $F^{(2)}_{\text{NN}}$, which may prevent the neural network from achieving sufficient learning, or from the training process neglecting the noise.
@@ 6 Conclusion
In this study, we proposed a novel method for estimating interactions among individuals within models of collective motion. This method combines dynamic GNNs with neural ODEs to estimate interactions among individuals. We demonstrated that this method was able to estimate interactions in both simple models and complex mixed-species collective motion models. In models with simple harmonic interactions, the proposed method identified the relevant interactions effectively. Importantly, the method successfully inferred the non-reciprocal interactions between the different species in the complex mixed-species model for the collective dynamics of slime molds. For the latter case, the continuous updating of the edge structures in our approach substantially reduces memory needs. In our simulations of 400 bodies with an edge density of 2%, the required memory of 200 GB for a fully connected graph is reduced to 30GB thus making them feasible on off-the-shelf GPUs. Nevertheless, we should note that this approach is limited due to the long time required for estimation. Moreover, the method currently estimates deterministic motion equations and only considers pairwise interactions, thereby not accommodating interactions among three or more bodies or the effects of noise. Future research is expected to extend this approach to estimate more general interactions and develop methods for stochastic motion equations. Applying the present methods to real data in systems such as immune cells should help clarify the complex rules behind their migration stategies.
@@ Acknowledgements
This research was supported by JSPS KAKENHI Grants JP22H05673 (awarded to MU), JP22H04841 (SKS), JP22K14012 (SKS), JP19H05799 (TJK), JP19H05801 (SS) and JP19KK0282 (SS), as well as JST CREST Grants JPMJCR2011 (TJK) and JPMJCR1923 (SS). We acknowledge the use of ChatGPT and GitHub Copilot for their assistance in swiftly articulating ideas in clear English. Passages written with the help of these tools were reviewed and rewritten by the authors.